# HWC-Loco: A Hierarchical Whole-Body Control Approach to Robust Humanoid Locomotion

### v2

@@ 4 Hierarchical Whole-Body Control for Humanoid Robust Locomotion
In practice, developing an end-to-end solver for the robust locomotion problem (4) is challenging, as it involves multiple deployment environments and constraints that must be adhered to. To tackle this issue, we propose dividing the objective into two stages: goal-tracking and safety recovery. The goal-tracking policy maximizes rewards while mimicking human behaviors (see Section 4.1). When mismatched environmental dynamics pose a risk to the safety of the control policy, HWC-Loco switches to the safety recovery policy, which manages safety-critical events to ensure the feasibility of humanoid control (see Section 4.2). To dynamically identify the ”sweet spot” for policy switching or activation, we introduce a high-level policy that coordinates the low-level policies by utilizing historical observations and the robot’s current status (see Section 4.3). Figure 1 illustrates the training pipeline of our HWC-Loco.
@@ 4.1 Learning Goal-Tracking Policy for Efficient Humanoid Locomotion
$\max_{\pi_{1}}\mathbb{E}_{\rho_{{M}^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{\mathcal{T}}^{L}}}}^{\pi_{1}}}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{\mathfrak{T}}(s_{t},a_{t})\right]{s.t.}\;\mathcal{D}_{f}(\rho_{{M}^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{\mathcal{T}}^{L}}}}^{\pi^{E}}\|\rho_{{M}^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{\mathcal{T}}^{L}}}}^{\pi_{1}})\leq\epsilon_{f}$
@@
Note that compared to the complete objective (4), this goal-tracking objective focuses only on maximizing performance in the training environments’ dynamics $P_{\mathcal{T}}^{L}$. Within this objective, the scale of the cumulative task rewards, $\sum_{t=0}^{\infty}\gamma^{t}r_{\mathfrak{T}}$ reflects how effectively the policy $\pi$ tracks locomotion-related commands, such as the robot’s target linear and angular velocities. Appendix A.2 specifies the definition of our task reward.
@@
In addition to the task rewards, a crucial aspect of the goal-tracking objective (LABEL:obj:goal-tracking) is evaluating the distance between the occupancy measures of the trained policy and the expert policy In the following, striving for concise representation, we represent $\rho_{{M}^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{\mathcal{T}}^{L}}}}^{\pi^{E}}$ and $\rho_{{M}^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}P_{\mathcal{T}}^{L}}}}^{\pi}$ by $\rho^{\pi^{E}}$ and $\rho^{\pi}$. Driven by information theory, previous works (Hussein et al., 2017) primarily implement $\mathcal{D}_{f}$ using metrics such as Kullback-Leibler (KL) divergence or Jensen-Shannon (JS) divergence. However, these metrics fail when the support of $\rho^{\pi^{E}}$ and $\rho^{\pi}$ has limited or zero overlap. In this study, we consider implement $\mathcal{D}_{f}$ with Wasserstein-1 distance under the Kantorovich-Rubinstein duality (Villani et al., 2009) such that:
$\mathcal{D}_{f}(\rho^{\pi^{E}}\|\rho^{\pi_{1}})=\sup_{\|f_{d}\|_{L}\leq 1}\mathbb{E}_{x\sim\rho^{\pi^{E}}}[f_{d}(x)]-\mathbb{E}_{x\sim\rho^{\pi_{1}}}[f_{d}(x)]$
@@
where the bounded Lipschitz-norm $\|f_{d}\|_{L}\leq 1$ ensure the smoothness of functions $f_{d}$. Inspired by adversarial imitation learning (Ho & Ermon, 2016; Zhang et al., 2020), an effective and intuitive implementation of $f_{d}$ is a discriminator for measuring whether $x$ is generated by human experts, such that $f_{d}:\mathcal{X}\rightarrow\{0,1\}$. This discriminator can be learned by maximizing the following objective:
$\max_{f_{d}}$ $\mathbb{E}_{x\sim\rho^{\pi^{E}}}[f_{d}(x)]-\mathbb{E}_{x\sim\rho^{\pi_{1}}}[f_{d}(x)]+$
$\beta\cdot\mathbb{E}_{\hat{x}\sim[\rho^{\pi^{E}},\rho^{\pi_{1}}]}\left[(\|\nabla_{\hat{x}}f_{d}(\hat{x})\|_{2}-1)^{2}\right]$ (7)
@@
where weighted by $\beta$, the gradient penalty term enforces the Lipschitz continuity required by the Wasserstein distance (i.e., $\|f_{d}\|_{L}\leq 1$). During training, $\rho^{\pi^{E}}$ represents the density of expert demonstration dataset $\mathcal{D}^{E}$, which is built from the CMU MoCap dataset (cmu, ). Additionally, we include data about humans standing, walking and running, enabling the robot to learn various human behaviors. We then retarget the motion data into a humanoid robot-compatible format, which we use as the training dataset.
Based on the discriminator, we can simplify the goal-tracking objective (LABEL:obj:goal-tracking) by substituting $\mathcal{D}_{f}$ with equation (LABEL:eqn:f-divergence):
@@
$\max_{\pi_{1}}\mathbb{E}_{\rho^{\pi_{1}}}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{\mathfrak{T}}(s_{t},a_{t})\right]$
${s.t.}\;\mathbb{E}_{\rho^{\pi^{E}}}[f_{d}(s^{d})]-\mathbb{E}_{\rho^{\pi_{1}}}[f_{d}(s^{d})]\leq\epsilon_{f}$ (8)
where $s^{d}$ is the input of the discriminator which is slight different from state $s$. Details can be seen in Appendix A.6.
@@
Given that this format of the constrained RL problem has a zero duality gap (Paternain et al., 2019), we can transform it into an unconstrained objective by analyzing the Lagrange dual form of the original constrained RL problem:
@@
where $\lambda$ denotes the optimal Lagrange multiplier. In practical implementation, the policy $\pi$ and discriminator $f_{d}$ can intuitively represent the generator and the discriminator under the adversarial learning framework. In implementation, we utilize the proximal policy optimization algorithm (Schulman et al., 2017) to update the goal-tracking policy. By alternatively updating (7) and (9), we can develop a goal-tracking policy that effectively maximizes task rewards while closely mimicking the expert’s behavior.
@@ 4.2 Learning Safety Recovery Policy for Handling Safe-Critical Events
Our safety recovery policy primarily focuses on handling emergency events, thereby recovering the robot from safety-critical situations and preventing failures of locomotion tasks, such as loss of balance or control policy malfunctions. To learn the safety-recover policy, we mainly study the complete objective (4). Similarly, as it is introduced in Section 4.1, we implement the divergence metric $\mathcal{D}_{f}$ with Wasserstein-1 distance, deriving the following objective:
@@
$\max_{\pi_{2}}\min_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\widehat{P_{\mathcal{T}}}}\in\mathfrak{P}_{\alpha}^{L}}\mathbb{E}_{\mu_{0},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\widehat{P_{\mathcal{T}}}},\pi_{2}}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{\mathfrak{T}}(s_{t},a_{t})-\lambda f_{d}(s_{t},a_{t})\right]$
${s.t.}\;\mathbb{E}_{\tau\sim(\mu_{0},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\widehat{P_{\mathcal{T}}}},\pi_{2})}\left[\phi(\tau)\right]\geq\epsilon_{\phi}$ (10)
@@
where $\epsilon_{\phi}$ is a hyper-parameter. By comparing this objective with the goal-tracking objective (9), both consider task rewards and human mimicking performance. Despite the similarity, key differences lie in the incorporation of uncertainty in environmental dynamics and feasibility constraints in policy updates. These additions necessitate the construction of an extreme-case uncertainty set and the modeling of safety constraints by learning the feasibility function $\phi(\cdot)$.
@@
Extreme-case Uncertainty Set. As in real-world deployment, there may have some complex terrains, external disturbances, hardware malfunctions, sensor noise, and errors caused by battery energy consumption. Those indicators may lead to extreme cases like falling and unexpected behavior. To model real-world extreme scenarios in the simulation (i.e., modelling ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\widehat{P_{\mathcal{T}}}}$s), We: 1) apply multi-scale external forces and torques to the humanoid body, 2) introduce random, high-intensity noise to the proprioceptive information and PD gains, 3) resample velocity commands to simulate malicious velocity inputs (Shi et al., 2024) within the goal-tracking policy, and 4) apply domain randomization to the training environment. Together, these dynamic variations construct the uncertainty set $\mathfrak{P}_{\alpha}^{L}$, under which we learn the safety-recover policy. Details of our implementation are provided in Appendix A.4.
@@
Zero-Moment Point (ZMP) Constraint. In the research of model-based robot control, a bipedal robot can be simplified to a linear inverted pendulum (Morisawa et al., 2006; Harada et al., 2006), where ZMP refers to the point where the ground reaction force has no horizontal moment. If the ZMP exits the support polygon (typically the foot’s contact area), the robot will quickly lose balance (Wieber, 2006; Feng et al., 2016; Scianca et al., 2020). Assuming a rigid body, stability is governed by gravity and the Center of Mass (CoM) accelerations, which also account for disturbances like obstacles and slippery surfaces. Under this assumption, ZMP is computed as follows:
$\mathbf{p}_{\text{ZMP}}(s,a)=\mathbf{p}_{\text{CoM}}(s,a)-\frac{z_{\text{CoM}}(s,a)}{g}\cdot\ddot{\mathbf{p}}_{\text{CoM}}(s,a)$ (11)
where 1) the vector $\mathbf{p}_{\text{CoM}}$ denotes the $x$ and $y$-coordinates of CoM position. 2) $z_{\text{CoM}}$ is the height of the CoM. 3) $g$ represents the acceleration due to gravity. 4) $\ddot{\mathbf{p}}_{\text{CoM}}$ indicates the accelerations of the CoM in the $x$ and $y$-directions, calculated from the current and previous velocities. By leveraging the property of ZMP constraints, we implement the feasibility indicator in the safety recovery objective (10) as follows:
$\phi(s,a)=\|\mathbf{p}_{\text{ZMP}}(s,a)-\mathbf{p_{\text{ac}}}\|_{2}$ (12)
@@
where $\mathbf{p}_{\text{ac}}$ represents the center of the support polygon (specific to each robot and known in advance). Intuitively, $\phi$ provides real-time insights into the robot’s stability and varies in its representation depending on the current support phase. Therefore, the humanoid robot can assess its stability in real time and use whole-body coordination to satisfy the ZMP constraint.
@@ 4.3 High-Level Planning for Policy Transition
The aforementioned low-level policies can achieve varying levels of optimality. A fundamental challenge is coordinating these policies to ensure that a humanoid robot can efficiently complete tasks while adhering to safety and robustness requirements. To address this challenge, we introduce a high-level planning policy that dynamically determines which policy to activate based on the specific state of humanoid robots. Considering the task of policy selection has a discrete action space, we utilize Double-DQN (Van Hasselt et al., 2016) to learn this high-level policy $\pi_{0}$ based on the following objective:
@@
where 1) the policy selection action $\bar{a}_{t}\in\{0,1\}^{2}$ denotes a one-hot vector featuring whether the goal-tracking policy $\pi_{1}$ and safety-recovery policy $\pi_{2}$, 2) $\mathds{1}(\bar{a}_{t-1}\neq\bar{a}_{\text{t}})$ denotes the continuality identifier for discouraging the over-frequent switch of policies, and 3) $\mathds{1}(s_{t})$ denotes the termination identifier for preventing the failing of locomotion tasks. (2) and (3) are both $r_{\mathrm{P}}$ defined in Section 3. More importantly, within this objective, by adjusting $\alpha$, practitioners can adjust the trade-off safety guarantee and task completion. Details of training process and Algorithm Pseudocode can be found in Appendix A.2.
Practical Deployment. The training of humanoid control policy commonly relies on the privileged information $\mathbb{P}_{t}$, such as external disturbances, terrain dynamics, etc. We follow  (Nahrendra et al., 2023) and infer $\mathbb{P}_{t}$ by training a VAE-based estimator $P(e_{t},z_{t}|o^{H}_{t})$ where $z_{t}$ captures the contextual embedding of the $\mathbb{P}_{t}$, and $e_{t}$ consists of body velocity and ZMP features. To enhance the representation of $\phi(\cdot)$, we introduce frequency encoding, similar to the approach in (Mildenhall et al., 2021), to capture the subtle variations in $\phi(\cdot)$ for the policy. Additionally, instead of the zero-mean Gaussian prior, we introduce a learnable prior similar to (Luo et al., 2023). This learnable prior allows the encoder to dynamically adjust the distribution based on the observations, facilitating a more effective representation of environmental features in the latent space. Moreover, we include a certain degree of domain randomization in the learning of HWC-Loco policies. The details of domain randomization are in Appendix A.4.

### v3

@@ 4 Hierarchical Whole-Body Control for Humanoid Robust Locomotion
In practice, developing an end-to-end solver for the robust locomotion problem (LABEL:obj:robust-crl) is challenging, as it involves multiple deployment environments and constraints that must be adhered to. To tackle this issue, we propose dividing the objective into two stages: goal-tracking and safety recovery. The goal-tracking policy maximizes rewards while mimicking human behaviors (see Section 4.1). When mismatched environmental dynamics pose a risk to the safety of the control policy, HWC-Loco switches to the safety recovery policy, which manages safety-critical events to ensure the feasibility of humanoid control (see Section 4.2). To dynamically identify the "sweet spot" for policy switching or activation, we introduce a high-level policy that coordinates the low-level policies based on historical observations and the robot’s current status (see Section 4.3). Figure 2 illustrates the training pipeline of our HWC-Loco.
@@ 4.1 Learning Goal-Tracking Policy for Efficient Humanoid Locomotion
$\max_{\pi_{1}}\mathbb{E}_{\rho_{M^{{\color[rgb]{1,0,0}P_{\mathcal{T}}^{L}}}}^{\pi_{1}}}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{\mathfrak{T}}(s_{t},a_{t})\right]{s.t.}\;\mathcal{D}_{f}(\rho_{M^{{\color[rgb]{1,0,0}P_{\mathcal{T}}^{L}}}}^{\pi^{E}}\|\rho_{M^{{\color[rgb]{1,0,0}P_{\mathcal{T}}^{L}}}}^{\pi_{1}})\leq\epsilon_{f}$ (5)
@@
Note that compared to the complete objective (LABEL:obj:robust-crl), this goal-tracking objective focuses only on maximizing performance in the training environments’ dynamics $P_{\mathcal{T}}^{L}$. Within this objective, the scale of the cumulative task rewards, $\sum_{t=0}^{\infty}\gamma^{t}r_{\mathfrak{T}}$ reflects how effectively the policy $\pi$ tracks locomotion-related commands. Appendix A.2 specifies the configuration of our task reward.
@@
In addition to the task rewards, a crucial aspect of the goal-tracking objective (5) is evaluating the distance between the occupancy measures of the trained policy and the expert policy In the following, striving for concise representation, we represent $\rho_{M^{{\color[rgb]{1,0,0}P_{\mathcal{T}}^{L}}}}^{\pi^{E}}$ and $\rho_{M^{{\color[rgb]{1,0,0}P_{\mathcal{T}}^{L}}}}^{\pi}$ by $\rho^{\pi^{E}}$ and $\rho^{\pi}$. Driven by information theory, previous works [47] primarily implement $\mathcal{D}_{f}$ using metrics such as Kullback-Leibler (KL) divergence or Jensen-Shannon (JS) divergence. However, these metrics fail when the support of $\rho^{\pi^{E}}$ and $\rho^{\pi}$ has limited or zero overlap. In this study, we consider implement $\mathcal{D}_{f}$ with Wasserstein-1 distance under the Kantorovich-Rubinstein duality [48] such that:
$\mathcal{D}_{f}(\rho^{\pi^{E}}\|\rho^{\pi_{1}})=\sup_{\|f_{d}\|_{L}\leq 1}\mathbb{E}_{x\sim\rho^{\pi^{E}}}[f_{d}(x)]-\mathbb{E}_{x\sim\rho^{\pi_{1}}}[f_{d}(x)]$ (6)
@@
where the bounded Lipschitz-norm $\|f_{d}\|_{L}\leq 1$ ensure the smoothness of functions $f_{d}$. Inspired by adversarial imitation learning [49, 50], an effective and intuitive implementation of $f_{d}$ is a discriminator for measuring whether $x$ is generated by human experts, such that $f_{d}:\mathcal{X}\rightarrow\mathbb{R}$. It can be learned by maximizing the following objective:
$\max_{f_{d}}\;\mathbb{E}_{x\sim\rho^{\pi^{E}}}[f_{d}(x)]-\mathbb{E}_{x\sim\rho^{\pi_{1}}}[f_{d}(x)]+\beta\cdot\mathbb{E}_{\hat{x}\sim[\rho^{\pi^{E}},\rho^{\pi_{1}}]}\left[(\|\nabla_{\hat{x}}f_{d}(\hat{x})\|_{2}-1)^{2}\right]$ (7)
@@
where weighted by $\beta$, the gradient penalty term enforces the Lipschitz continuity required by the Wasserstein distance (i.e., $\|f_{d}\|_{L}\leq 1$). During training, $\rho^{\pi^{E}}$ represents the density of expert demonstration dataset $\mathcal{D}^{E}$, which is built from the CMU MoCap dataset [51]. Additionally, we include data about humans standing, walking, and running to learn various human behaviors. We then retarget the motion data into a humanoid robot-compatible format, which we use as the training dataset. By substituting $\mathcal{D}_{f}$ with equation (6), the goal-tracking objective (5) can be simplified as:
@@
$\max_{\pi_{1}}\mathbb{E}_{\rho^{\pi_{1}}}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{\mathfrak{T}}(s_{t},a_{t})\right]\quad\text{s.t.}\quad\mathbb{E}_{\rho^{\pi^{E}}}[f_{d}(s^{d})]-\mathbb{E}_{\rho^{\pi_{1}}}[f_{d}(s^{d})]\leq\epsilon_{f}$ (8)
@@
where $s^{d}$ records the necessary state information for the discriminator. (See details in Appendix A.5). Given that this formulation of constrained RL problem has a zero duality gap [52], we can transform it into an unconstrained objective by analyzing the Lagrange dual form of the original constrained RL problem:
@@
where $\lambda$ denotes the optimal Lagrange multiplier. In practical implementation, the policy $\pi$ and discriminator $f_{d}$ can intuitively represent the generator and the discriminator under the adversarial learning framework. In implementation, we utilize the proximal policy optimization algorithm [53] to update the goal-tracking policy. By alternatively updating (7) and (9), we can develop a goal-tracking policy that effectively maximizes task rewards while closely mimicking the expert’s behavior.
@@ 4.2 Learning Safety Recovery Policy for Handling Safe-Critical Events
Our safety recovery policy primarily focuses on handling emergency events, thereby recovering the robot from safety-critical situations and preventing failures of locomotion tasks, such as loss of balance or control policy malfunctions. To learn the safety-recover policy, we mainly study the complete objective (LABEL:obj:robust-crl). Similarly, as it is introduced in Section 4.1, we implement the divergence metric $\mathcal{D}_{f}$ with Wasserstein-1 distance, deriving the following objective:
@@
$\max_{\pi_{2}}\min_{{\color[rgb]{0,0,1}\widehat{P_{\mathcal{T}}}}\in\mathfrak{P}_{\alpha}^{L}}\mathbb{E}_{\mu_{0},{\color[rgb]{0,0,1}\widehat{P_{\mathcal{T}}}},\pi_{2}}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{\mathfrak{T}}(s_{t},a_{t})-\lambda f_{d}(s_{t},a_{t})\right]\;\;{s.t.}\;\;\mathbb{E}_{\tau\sim(\mu_{0},{\color[rgb]{0,0,1}\widehat{P_{\mathcal{T}}}},\pi_{2})}\left[\phi(\tau)\right]\geq\epsilon_{\phi}$ (10)
@@
where $\epsilon_{\phi}$ is a hyperparameter. By comparing this objective with the goal-tracking objective (9), while both consider task rewards and human mimicking performance, their key differences lie in the incorporation of uncertainty in environmental dynamics and feasibility constraints in policy updates. These additions necessitate the construction of an extreme-case uncertainty set and the modeling of safety constraints by learning the feasibility function $\phi(\cdot)$.
@@
Extreme-case Uncertainty Set. Real-world deployment has complex terrains, external disturbances, hardware malfunctions, and sensor noise. Those factors may lead to extreme cases like falling or unexpected behaviors. To model them in the simulation (i.e., modelling ${\color[rgb]{0,0,1}\widehat{P_{\mathcal{T}}}}$s), We: 1) apply multi-scale external forces and torques to the humanoid body, 2) introduce random, high-intensity noise to the proprioceptive information and PD gains, 3) resample velocity commands to simulate malicious velocity inputs [54] within the goal-tracking policy, and 4) apply domain randomization to the training environment. Together, these dynamic variations construct the uncertainty set $\mathfrak{P}_{\alpha}^{L}$, under which we learn the safety-recover policy. Appendix A.7 shows our implementation details.
@@
Zero-Moment Point (ZMP) Constraint. A bipedal robot can be generally modeled by a linear inverted pendulum [55, 56], where ZMP refers to the point where the ground reaction force has no horizontal moment. If the ZMP exits the support polygon (typically the foot’s contact area), the robot will quickly lose balance [57, 58, 59, 60]. Stability is governed by gravity and the Center of Mass (CoM) accelerations, which also account for disturbances like obstacles and slippery surfaces. Under this assumption, we implement the feasibility indicator in the safety recovery objective (10) as follows:
$\phi(s,a)=\|\mathbf{p}_{\text{ZMP}}(s,a)-\mathbf{p_{\text{ac}}}\|_{2}\;\;\text{where}\;\;\mathbf{p}_{\text{ZMP}}(s,a)=\mathbf{p}_{\text{CoM}}(s,a)-\frac{z_{\text{CoM}}(s,a)}{g}\cdot\ddot{\mathbf{p}}_{\text{CoM}}(s,a)$ (11)
@@
where 1) the vector $\mathbf{p}_{\text{CoM}}$ denotes the $x$ and $y$-coordinates of CoM position. 2) $z_{\text{CoM}}$ is the height of the CoM. 3) $g$ represents the acceleration due to gravity. 4) $\ddot{\mathbf{p}}_{\text{CoM}}$ indicates the accelerations of the CoM in the $x$ and $y$-directions. 5) $\mathbf{p}_{\text{ac}}$ represents the center of the support polygon (specific to each robot and known in advance). Intuitively, $\phi$ provides real-time insights into the robot’s stability and varies in its representation depending on the current support phase. Therefore, the humanoid robot can assess its stability in real time and use whole-body coordination to satisfy the ZMP constraint.
@@ 4.3 High-Level Planning for Policy Transition
Figure 3: Proportion of goal-tracking policy.
While the low-level policies can achieve varying levels of optimality, a fundamental challenge is coordinating these policies to ensure that a humanoid robot can efficiently complete tasks while adhering to safety and robustness requirements. We introduce a high-level planning policy that dynamically determines which policy to activate based on the specific state of humanoid robots. The task of policy selection has a discrete action space, and we learn this high-level policy $\pi_{0}$ based on the following objective:
@@
where 1) the policy selection action $\bar{a}_{t}\in\{0,1\}^{2}$ denotes a one-hot vector featuring whether the goal-tracking policy $\pi_{1}$ and safety-recovery policy $\pi_{2}$, 2) $\mathds{1}(\bar{a}_{t-1}\neq\bar{a}_{\text{t}})$ denotes the continuality identifier for discouraging the over-frequent switch of policies, and 3) $\mathds{1}(s_{t})$ denotes the termination identifier for preventing the failing of locomotion tasks. (2) and (3) are both $r_{\mathrm{P}}$ defined in Section 3. More importantly, within this objective, by adjusting $\alpha$, practitioners can adjust the trade-off safety guarantee and task completion. Training and algorithmic details can be found in Appendix A.2.
We evaluate the average activation proportion of the goal-tracking policy during training. As shown in Figure 3, when $\alpha=0$, 20, or 50, the goal-tracking policy $\pi_{1}$ is predominantly selected, with the recovery policy $\pi_{2}$ triggered only in rare dangerous cases. In contrast, when $\alpha=200$, the sparsity of termination rewards induces training instability, resulting in overly conservative behavior.
Practical Deployment. The training of humanoid control policy commonly relies on the privileged information $\mathbb{P}_{t}$, such as external disturbances, terrain dynamics, etc. We follow  [18] and infer $\mathbb{P}_{t}$ by training a VAE-based estimator $P(e_{t},z_{t}|o^{H}_{t})$ where $z_{t}$ captures the contextual embedding of the $\mathbb{P}_{t}$, and $e_{t}$ consists of body velocity and ZMP features. To enhance the representation of $\phi(\cdot)$, we introduce frequency encoding, similar to the approach in [61], to capture the subtle variations in $\phi(\cdot)$ for the policy. Additionally, instead of the zero-mean Gaussian prior, we introduce a learnable prior similar to [62]. Moreover, we include a certain degree of domain randomization in the learning of HWC-Loco policies. The details of domain randomization are in Appendix A.7.